# Poker Hand History File Format Specification

### v2

@@ Abstract
This paper introduces the Poker Hand History (PHH) file format, designed to standardize the recording of poker hands across different game variants. Despite poker’s widespread popularity in the mainstream culture as a mind sport and its prominence in the field of artificial intelligence (AI) research as a benchmark for imperfect information AI agents, it lacks a consistent format that humans can use to document poker hands across different variants that can also easily be parsed by machines. To address this gap in the literature, we propose the PHH format which provides a concise human-readable machine-friendly representation of hand history that comprehensively captures various details of the hand, ranging from initial game parameters and actions to contextual parameters including but not limited to the venue, players, and time control information. In the supplementary, we provide over 10,000 hands covering 11 different variants in the PHH format. Building on our previous work on PokerKit, a premier poker hand simulation tool, we demonstrate the usages of our open-source Python implementation of the PHH parser. The source code of the parser is available on GitHub: https://github.com/uoftcprg/pokerkit
@@ I Introduction
Poker is a popular card game with a stochastic nature and imperfect information aspect that has not only captured the mainstream attention as a leading mind sport but also served as a benchmark of artificial intelligence (AI) research, namely in the development of imperfect information AI agents.
Pioneering breakthroughs in imperfect information game-solving algorithms have often been accompanied by their applications in state-of-the-art poker AI agents as seen in the example of Rhode Island hold ’em [1], fixed-limit Texas hold ’em [2], heads-up no-limit Texas hold ’em [3, 4], and 6-player no-limit Texas hold ’em [5]. Successive state-of-the-art algorithms were shown to solve games of progressively larger game tree sizes.
While impressive, the applications of these algorithms were mostly limited to the Texas hold ’em variant. There exists countless variants of poker that introduce different rules and actions not present in Texas hold ’em [6, 7] that can serve as new challenges in the development of imperfect information AI agents. For instance, Omaha variants involve four hole cards per player instead of simply two, which increases the range of possible holdings of opponents, further increasing the game tree size. Another popular variant, lowball draw, introduces shifting hidden information depending on the discards and draws unlike Texas hold ’em where the players’ hole cards remain unchanged throughout the hand.
@@
However, our previous efforts in the digitalization of poker were only partially complete, as we provided no means to record poker hands for persistent storage and subsequent analysis. This is partly due to the fact that there does not exist a standardized format for recording poker games that is not only human-friendly but also easily parsable by computer software akin to the portable game notation (PGN) for chess [9] or smart game format (SGF) for go [10].
This paper proposes a novel open Poker Hand History (PHH) file format designed to provide a consistent method to annotate poker hands across different variants in addition to capturing other details of the game including but not limited to the venue, players, and time control information that can aid in domains such as poker AI agent development, historical analysis, and data referencing.
In the next section, we provide some motivational examples that showcase the versatility of the PHH notation. This is followed by an exploration of the related works in the literature that introduce alternatives to our proposal with which we later run benchmarks for comparison purposes. Then, we provide the detailed specification of the PHH format, listing the required and optional fields used to comprehensively describe poker hands in addition to various recommendations in file formatting, style guides, and parser implementation. Finally, for demonstrative purposes, we provide sample usages of our open-source Python PHH reader and writer which uses PokerKit as the simulation backend.
In the supplementary, over 10,000 sample poker hands, covering 11 different poker variants and composed of small selections of historical or notable hands, all 83 televised hands in the 2023 WSOP Event #43: \$50,000 Poker Players Championship (PPC) — Day 5 [11], and the hands played by Pluribus in Brown and Sandholm [5], are provided in the PHH file format.
II Motivational Examples
# A big pot between Dwan and Ivey.
variant = "NT"
antes = [500, 500, 500]
blinds_or_straddles = [1000, 2000, 0]
min_bet = 2000
starting_stacks = [
  1125600,
  2000000,
  553500,
actions = [
  # Pre-flop
  "d dh p1 Ac2d",  # Ivey
  "d dh p2 ????",  # Antonius
  "d dh p3 7h6h",  # Dwan
  "p3 cbr 7000",  # Dwan
  "p1 cbr 23000",  # Ivey
  "p2 f",  # Antonius
  "p3 cc",  # Dwan
  # Flop
  "d db Jc3d5c",
  "p1 cbr 35000",  # Ivey
  "p3 cc",  # Dwan
  # Turn
  "d db 4h",
  "p1 cbr 90000",  # Ivey
  "p3 cbr 232600",  # Dwan
  "p1 cbr 1067100",  # Ivey
  "p3 cc",  # Dwan
  "p1 sm Ac2d",  # Ivey
  "p3 sm 7h6h",  # Dwan
  # River
  "d db Jh",
author = "Juho Kim"
event = "Million Dollar Cash Game"
year = 2009
players = [
  "Phil Ivey",
  "Patrik Antonius",
  "Tom Dwan",
currency = "USD"
Figure 1: The first televised million-dollar pot between Tom Dwan and Phil Ivey [12] in the PHH file format. Note that the starting stack of Patrik Antonius is unknown.
@@ II Motivational Examples
Figures 1 and 2 show two historical poker hands digitalized into the PHH format. The former involves the first televised million-dollar pot between Tom Dwan and Phil Ivey in the 2009 Full Tilt Million Dollar Cash Game Season 4 Episode 12 [12]. The latter describes a bad beat from the final table of the 2019 WSOP Event #52: \$50,000 PPC on Day 5 of the event between Josh Arieh and Bryce Yockey [13].
@@ III Related Works
In poker literature, it has historically been the norm to describe poker games verbally as exemplified in the news articles covering the aforementioned hands [12, 13]. Such a practice, while fit for human readership, presents significant challenges for parsing and computational analysis. While the description could be rich with information like the names of players involved, the backstory of the hand, et cetera, it could also lack nuances and details that may be of interest to the readers such as the exact stack sizes, bet sizes, and mucked hand values, just to name a few. The subjective formatting may introduce bias and inconsistencies unsuitable for scientific use cases.
@@
PokerStars Hand #????????????: Holdem No Limit (\$1000.00/\$2000.00) - 2009/??/?? ??:??:?? ??? [2009/??/?? ??:??:?? ET]
Table ‘?????’ ?-max Seat #3 is the button
Seat 1: PhilIvey (\$1125600.00 in chips)
Seat 2: PatrikAntonius (\$2000000.00 in chips)
Seat 3: Hero (\$553500.00 in chips)
PhilIvey: posts the ante \$500.00
PatrikAntonius: posts the ante \$500.00
Hero: posts the ante \$500.00
PhilIvey: posts small blind \$1000.00
PatrikAntonius: posts big blind \$2000.00
*** HOLE CARDS ***
Dealt to Hero [7h 6h]
Hero: raises \$5000.00 to \$7000.00
PhilIvey: raises \$16000.00 to \$23000.00
PatrikAntonius: folds
Hero: calls \$16000.00
*** FLOP *** [Jc 3d 5c]
PhilIvey: bets \$35000.00
Hero: calls \$35000.00
*** TURN *** [Jc 3d 5c] [4h]
PhilIvey: bets \$90000.00
Hero: raises \$142600.00 to \$232600.00
PhilIvey: raises \$834500.00 to \$1067100.00 and is all-in
Hero: calls \$262400.00
Uncalled bet (\$572100.00) returned to PhilIvey
*** RIVER *** [Jc 3d 5c 4h] [Jh]
*** SHOW DOWN ***
PhilIvey: shows [Ac 2d] (a straight, Ace to Five)
Hero: shows [7h 6h] (a straight, Three to Seven)
Hero collected \$1109500.00 from pot
*** SUMMARY ***
Total pot \$1109500.00 | Rake ?
Board [Jc 3d 5c 4h Jh]
Figure 3: The identical hand played in Figure 1 converted to the PokerStars hand history format. Note that unknown information is marked with question marks and the first line was broken to fit in the page. Again, the starting stack of Patrik Antonius is unknown.
PokerStars, one of the most popular online poker platforms, saves a log of the game history in their PokerStars Hand History format [15]. As an example, the identical hand shown in Figure 1, converted into the PokerStars hand history format, is shown in Figure 3. While the logs are designed for human readership and record-keeping, they are in plaintext format and are highly structured and consistent, allowing various poker analytic tools in the industry such as Holdem Manager and Poker Copilot to parse and analyze a player’s hand history reliably. The logged details include granular information down to the amounts of antes and blinds being posted by each player. This fine detail unfortunately makes it too verbose to be hand-typed by a human while being too specialized for use in video poker. For instance, some digital archiving use cases may require additional or complete descriptions of the game like the final stacks, time control, et cetera. Not only that, the verbal nature of the format introduces unnecessary verbosity that wastes storage space.
STATE:7:fr225fffc/cr475c/cr1225c/cc:5hJc|Jd9h|6s5c|Ah7h|2s2d|3hTs/3sJh2h/Tc/Ks:-50|1275|0|-1225|0|0:Gogo|Budd|Eddie|Bill|Pluribus|MrWhite
Figure 4: A sample hand played by Pluribus in the supplementary of Brown and Sandholm [5]. Note that the line was broken to fit in the page.
The dataset of hands played by Pluribus, in the supplementary of Brown and Sandholm [5], is formatted in a much more concise way as shown in Figure 4. This format includes the bare minimum description of actions to comprehensively represent a no-limit Texas hold ’em hand. However, due to the fact that each action notation does not specify the actor, the hands are difficult to follow by a human reader. In addition, the format is limited to only a single variant: no-limit Texas hold ’em, and makes critical assumptions such as the identical and fixed starting stack sizes of 100 big blinds and the number of players being 6. On top of this, it lacks options to include various metadata outside of the game state descriptions that may be of interest for data referencing purposes.
@@ IV Objective
An ideal representation of poker hands should not only be concise but also be easily readable and writable by human readers. It should also support only describing the bare minimum information necessary to comprehensively describe a hand while facilitating options to give rich details for various use cases beyond simple action tracking. For convenience in parser implementations, the formatting must also be simply structured in a consistent manner. In addition, the format must be powerful enough to describe a wide variety of poker games played in formal settings. Such selections should cover games played in reputable poker tournaments around the world while the inclusion of various eccentric or esoteric variants is not as necessary.
@@ V Specification
The PHH format is a derivative of the Tom’s Obvious, Minimal Language (TOML) format [16]. As such, all PHH files must also be valid TOML files, following the TOML syntax. This design allows PHH files to be human-readable and writable while maintaining easy parsability by software systems. Software parser implementations are encouraged to utilize built-in or third-party tools to read and write TOML files.
The PHH format puts restrictions on the naming and types of the key/values, with which the poker games are described. These definitions, named “fields,” fall under the following three classifications:
  1. 1.
    Initial state construction.
  2. 2.
    Subsequent state progression.
  3. 3.
    Miscellaneous game information.
The first and second classifications, involving state construction and progression, are grouped together as “required” fields and must be specified. The third classification is dubbed as “optional” fields and may be omitted by the annotator.
@@ V-B Position
In typical button games such as no-limit Texas hold ’em or pot-limit Omaha hold ’em, the first player must be in the small blind while the last player must be in position (i.e. has the button). In non-button games like stud, there is no strict notion of position. In such games, the first player should be the one in the dealer’s immediate left while the last player should be in the dealer’s immediate right.
@@ V-C Rules
As poker is a game with a diverse set of rules that often differ based on regions or events, it is impossible for us to point to anything as an authoritative documentation of poker rules that must be followed both by the annotator and the parser. The parser implementation should be designed to handle actions that may be potential transgressions in certain poker rule sets, either silently or with a warning.
@@ VI Required Fields
Variant variant String
@@
Blinds or Straddles blinds_or_straddles Array of integers or floats
@@
Starting stacks starting_stacks Array of integers or floats
@@
The required fields represent a handful of fields that the annotator must define to digitalize a poker hand and therefore form a valid PHH file. The required fields are either used during the initial state construction or during the subsequent state progression.
In this paper and the supplementary, 11 variants are explored, each of which is written in an abbreviated form, as tabulated in Table I. These variants include all 9 played in the 2023 WSOP Event 43: \$50,000 PPC [11], in addition to 2 others like short-deck and badugi in which the poker community has shown growing interest. These were chosen as they adequately represent the diverse nature of poker variants and demonstrate the flexibility of the specification laid out in this paper which is powerful enough to accommodate all variants in the 2023 World Series of Poker Tournament Rules [6]. Those not mentioned in Table I can simply be integrated by introducing new variant codes to be associated with, and they will be introduced on a need-by-need basis based on community feedback in the future version of the specification.
VI-A State Construction
The fields in this section concern itself with the topic of the initial state construction.
VI-A1 Variant
The “variant” field, which has a string value, defines what variant the hand is in and therefore determines the composition of the required fields that must be specified. The field value is a code unique to each variant, 11 of which are tabulated in Table I, along with the corresponding field requirements. Sample definitions are shown in Figures 1 and 2.
VI-A2 Antes
The value of the “antes” field, representing the antes, must be an array of non-negative integers or floats of length equal to the number of players. In general, there are two types of antes: uniform antes and non-uniform antes. In the case of uniform antes, all the array elements are identical, meaning the antes are the same for each player. No antes (zero antes) must be represented as an array of zero integers or floats while non-zero uniform antes can be represented as an array of constant positive TOML native numerical values. Non-uniform ante formats such as big-blind antes and button antes must be represented as an array with non-zero values in the indices corresponding to the anted players. Example definitions are shown in Figure 5.
@@ VI-A2 Antes
Figure 5: Example ante definitions in PHH format.
VI-A3 Blinds or Straddles
The “blinds_or_straddles” field, representing all forced bets except the bring-in, must only be specified in button games as an array of non-negative integers or floats of length equal to the number of players. Most button games have two blinds: a small blind and a big blind. These are simply described by an array with two positive values in the first two indices while all other elements are zero. Optionally, various straddle regimes ranging from under-the-gun straddle, double straddle, button straddle, et cetera, can be introduced by setting the elements in the relevant indices with non-zero values. Examples are demonstrated in Figure 6. Note that, in heads-up situations, the effective blind amounts are the reversed version of the array field value.
@@ VI-A3 Blinds or Straddles
Figure 6: Example blind or straddle definitions in PHH format.
VI-A4 Bring-in
The “bring_in” field is of a positive integer or float value and is primarily used in stud poker variants to represent the bring-in forced bet. The usage of bring-ins is mutually exclusive with blinds or straddles. In other words, both must never be defined together. A sample definition is shown in Figure 7.
bring_in = 1
small_bet = 2
big_bet = 4
Figure 7: Selections of required field definition in PHH format.
VI-A5 Small bet
The “small_bet” field, a positive integer or a float, denotes the minimum bet in the first few betting rounds of fixed-limit games. It is not a feature of pot-limit or no-limit games. Sample use cases are shown in Figures 2 and 7.
VI-A6 Big bet
The “big_bet” field, a positive integer or a float, similar to its smaller counterpart, denotes the minimum bet in the last few betting rounds of fixed-limit games. Again, this field is not featured in unfixed-limit games. sample definitions are shown in Figures 2 and 7.
VI-A7 Min bet
The “min_bet” field, a positive integer or a float, denotes the minimum bet in all of the betting rounds of no-limit or pot-limit games. This field must not be specified in fixed-limit games. A sample usage is shown in Figure 1.
VI-A8 Starting stacks
The “starting_stacks” field denotes the starting stacks of each player and is an array of positive integers or floats of length equal to the number of players. All stack sizes must strictly be non-zero. Sample usages are shown in Figures 1 and 2.
VI-B State Progression
The one and only field that represents the progression of a poker state is the “actions” field. This field is an array of string action representations, following our action notation described below. It describes how the state mutates throughout the hand as the actors – the dealer and the players – act.
VI-B1 Actions
The “actions” field is an array of string action notations that describe what each actor did in their turn. For the dealer (an actor), their action notations denote what cards were dealt to each player or to the board. More specifically:
  • Dealing board cards or
  • Dealing hole cards.
For each player, their action notations may describe them:
  • Posting a bring-in;
  • Completing;
  • Betting;
  • Raising or raising to;
  • Checking;
  • Calling;
  • Folding;
  • Standing pat;
  • Discarding;
  • Showing their hole cards; or
  • Mucking their hole cards.
The notations in the “actions” field must be sorted in the order of the action. Its first element, if the array is not empty, must be the first action taken, which, in all mainstream poker variants, is the hole card dealing action by the dealer. The actions may represent a complete history of the hand or a partial history that does not reach the terminal state (the hand being over).
Figures 1 and 2, as well as the PHH files in the supplementary, demonstrate the sample usages of this field.
VI-C Action Notation
@@
Note that commentaries and TOML comments are two distinct mechanisms in the PHH format. They are both followed by a single hash symbol. However, TOML comments are ignored by the parser, while the commentary information is considered a part of the action notation by the parsing system. The annotator should use commentaries to give extra contextual information about the hand that cannot be determined by the parser such as a summary of the metagame between some of the players, exposed hole cards, and interruptions, just to name a few, whereas the TOML comments should serve as a guide to the reader who is directly reading the file without the help of a software parser. For instance, the information provided by TOML comments may include the names of the betting round or the actor.
@@
VI-C1 Board Dealing
[…]
is the board dealing action where the dealer deals the community cards to the board at the beginning of each applicable street. The notation of the board dealing action must include a single argument of a valid card notation representing the newly dealt community cards. The correct number of cards must be dealt; the cards must all be notated as known cards. Sample action notations are written in Figure 1.
VI-C2 Hole Dealing
The hole dealing action
[…]
represents the action of dealing hole cards to a particular player and must accept two arguments, the player being dealt and the newly dealt hole cards (in card notation) some or all of which may be unknown. Its usages are shown in Figures 1 and 2. As in the board dealing action, the correct number of hole cards must be dealt to each player. The hole dealing must be made starting from the first active player to the last active player still in the hand.
VI-C3 Standing Pat or Discarding
The standing pat and discarding actions
[…]
are grouped together due to their similarity. Standing pat is a special case of discarding where a player does not discard any hole cards. This notation accepts an optional argument denoting the cards that are being discarded. An omission of this argument denotes that the player in turn is standing pat while the inclusion denotes that the player discarding. All discarded cards must be part of the player’s hole cards. Therefore, the card notation in the argument may contain unknown cards if and only if some of the player’s hole cards are unknown. This action is demonstrated in Figure 2.
VI-C4 Bring-in Posting
This action, with the
$\text{p}n\text{ pb}[\text{ \# Commentary}]$
grammar, denotes the posting of the bring-in forced bet. A sample notation of such action is: “p3 pb” which describes that Player 3 is bringing-in.
VI-C5 Folding
$\text{p}n\text{ f}[\text{ \# Commentary}]$
is the folding action. Examples are shown in Figures 1 and 2.
VI-C6 Checking or Calling
This action represents checking or calling
$\text{p}n\text{ cc}[\text{ \# Commentary}]$
which entails matching the largest bet. When the amount required to match is zero, the player is said to check. Otherwise (in the case of a non-zero amount), the player is said to call. The examples are shown in Figures 1 and 2.
VI-C7 Completion, Betting, or Raising to
This action represents completion, betting, or raising to a positive integral or real amount, supplied as an argument, and uses the
$\text{p}n\text{ cbr amount}[\text{ \# Commentary}]$
grammar. This positive amount must be between the valid ranges set by the betting structure of the variant. When this action is performed as an alternative to or in the face of a bring-in, the player is said to complete. Alternatively, when nobody else bet, the player is said to be betting. Otherwise, the player is said to be making a raise. These are demonstrated in Figures 1 and 2.
VI-C8 Showing or Mucking Hole Cards
[…]
action is the only action that can be performed during showdown or all-in situations. It represents showing of the hole cards to try to win the pot or after an all-in, or mucking of the hole cards. This action accepts an optional argument of cards which, if supplied, must all be known. The omission of this field represents mucking while the inclusion represents showing. Instances of this action are shown in Figures 1 and 2.
VI-C9 No operations
No-ops
[…]
are represented with an empty string, a string composed entirely of whitespace characters like “ ”, or a standalone commentary like “# Burn card 6s is exposed.”
VI-D Card Notation
TABLE IV: All ranks and their single character representations.

### v3

# Recording and Describing Poker Hands

@@ Abstract
This paper introduces the Poker Hand History (PHH) file format, designed to standardize the recording of poker hands across different game variants. Despite poker’s widespread popularity in the mainstream culture as a mind sport and its prominence in the field of artificial intelligence (AI) research as a benchmark for imperfect information AI agents, it lacks a consistent format that humans can use to document poker hands across different variants that can also easily be parsed by machines. To address this gap in the literature, we propose the PHH format which provides a concise human-readable machine-friendly representation of hand history that comprehensively captures various details of the hand, ranging from initial game parameters and actions to contextual parameters including but not limited to the venue, players, and time control information. In the supplementary, we provide 10,088 hands covering 11 different variants in the PHH format. The source code of the parser is available on GitHub: https://github.com/uoftcprg/pokerkit
@@ I Introduction
Poker is a popular card game with a stochastic nature and imperfect information aspect that has not only captured mainstream attention as a leading mind sport but also served as a benchmark of artificial intelligence (AI) research, namely in the development of imperfect information AI agents.
Pioneering breakthroughs in imperfect information game-solving algorithms have often been accompanied by their applications in state-of-the-art poker AI agents as seen in the example of Rhode Island hold ’em [1], fixed-limit Texas hold ’em [2], heads-up (2-player) no-limit Texas hold ’em [3, 4], and 6-max (6-player) no-limit Texas hold ’em [5]. Successive state-of-the-art algorithms were shown to solve games of progressively larger game tree sizes.
While impressive, the applications of these algorithms were mostly limited to the Texas hold ’em variant which ignores the vast diversity of poker. Indeed, there exists countless variants of poker that introduce different rules and actions not present in Texas hold ’em [6, 7] that can serve as new challenges in the development of imperfect information AI agents. For instance, Omaha variants involve four hole cards per player instead of simply two, which increases the range of possible holdings of opponents, further increasing the game tree size. Another popular variant, lowball draw, introduces shifting hidden information depending on the discards and draws unlike Texas hold ’em where the players’ hole cards remain unchanged throughout the hand.
While it is true that Texas hold ’em is the most prominent variant of poker, as it is the variant of choice in the World Series of Poker (WSOP) Main Event, the largest poker tournament in the world, there is a substantial interest in the market and the industry on other poker variants. For instance, of the 94 in-person bracelet events in the 2023 WSOP, only around 51 events (54.26%) play no-limit Texas hold ’em exclusively. [8] This diversity of the game differs from other popular board games like chess, shogi, or go which has a universally accepted “standard” variant.
In addition, various initial state assumptions made by the state-of-the-art poker AI agents make it completely inapplicable to real-life gameplays where the assumptions are rarely true. All the recent breakthroughs in poker AI agents deal with uniform starting stack sizes, where every player has an identical number of chips at the beginning of the game, in addition to just one configuration of antes and forced bets like blinds or straddles. Not to mention, their solutions are only meant to be used in either heads-up or 6-max settings. These assumptions, of course, are unrealistic, and taking these variations into account vastly increases the game tree size. It is therefore uncertain if existing techniques for solving imperfect information games can handle game trees of such size. The implications of the existence of such an algorithm would be immense, given that the vast number of real-world problems such as stock markets, auctions, and diplomacy are examples of imperfect information games like poker.
@@
However, our previous efforts in the digitalization of poker were only partially complete, as we provided no means to record poker hands for persistent storage and subsequent analysis, a crucial component in the training and evaluation of AI agents. This is partly because there does not exist a standardized format for recording poker games that is not only human-friendly but also easily parsable by computer software akin to the portable game notation (PGN) for chess [10] or smart game format (SGF) for go [11].
This paper proposes a novel open-source Poker Hand History (PHH) file format designed to provide a consistent method to annotate poker hands across different variants in addition to capturing other details of the game including but not limited to the venue, players, and time control information that can aid in domains such as poker AI agent development, historical analysis, and data referencing.
In the next section, we provide a motivational example that showcase the versatility of the PHH notation. This is followed by an exploration of the related works in the literature that introduce alternatives to our proposal with which we later run benchmarks for comparison purposes. Then, we provide the detailed specification of the PHH format, listing the required and optional fields used to comprehensively describe poker hands in addition to various recommendations in file formatting, style guides, and parser implementation.
In the supplementary, 10,088 sample poker hands, covering 11 different poker variants and composed of small selections of historical or notable hands, all 83 televised hands in the 2023 WSOP Event #43: \$50,000 Poker Players Championship (PPC) — Day 5 [12], and the hands played by Pluribus in Brown and Sandholm [5], are provided in the PHH file format.
II Motivational Example
@@
Figure 1 shows an example of a poker hand digitalized into the PHH format from a bad beat from the final table of the 2019 WSOP Event #52: \$50,000 PPC on Day 5 of the event between Josh Arieh and Bryce Yockey [13].
@@ III Related Works
In poker literature, it has historically been the norm to describe poker games verbally as exemplified in the news articles covering the aforementioned hands [14, 13]. Such a practice, while fit for human readership, presents significant challenges for parsing and computational analysis. While the description could be rich with information like the names of players involved, the backstory of the hand, et cetera, it could also lack nuances and details that may be of interest to the readers such as the exact stack sizes, bet sizes, and mucked hand values. The subjective formatting may introduce bias and inconsistencies unsuitable for scientific use cases.
@@
PokerStars, one of the most popular online poker platforms, saves a log of the game history in their proprietary PokerStars Hand History format [16]. While the logs are designed for human readership and record-keeping, they are in plaintext format and are highly structured and consistent, allowing various poker analytic tools in the industry such as Holdem Manager, Poker Copilot, and PokerSnowie to parse and analyze a player’s hand history reliably. The logged details include granular information down to the amounts of antes and blinds being posted by each player. This fine detail unfortunately makes it too verbose to be hand-typed by a human while being too specialized for use in video poker, namely PokerStars. For instance, some digital archiving use cases may require additional or complete descriptions of the game like the final stacks, time control, et cetera with features of the game not available on PokerStars. Not only that, the verbal nature of the format introduces unnecessary verbosity that wastes storage space. This flaw is further exacerbated by its lack of sufficient specification or documentation which introduces substantial difficulties when developing datasets in the PokerStars format. This is evident by the fact that there has not been any notable community effort to digitalize any historical poker hands.
The dataset of hands played by Pluribus, in the supplementary of Brown and Sandholm [5], is formatted extremely concisely where a single hand is summarized into a dense line of characters. This format includes the bare minimum description of actions to comprehensively represent a no-limit Texas hold ’em hand assuming uniform starting stacks and standard double blinds. However, since each action notation does not specify the actor, the hands are difficult to follow by a human reader. In addition, the format is limited to only a single variant: no-limit Texas hold ’em, and makes critical assumptions such as the identical and fixed starting stack sizes of 100 big blinds and the number of players being 6. On top of this, it lacks options to include various metadata outside of the game state descriptions that may be of interest for data referencing purposes.
@@ IV Objective
An ideal representation of poker hands should not only be concise but also easily readable and writable by human readers. It should also support only describing the bare minimum information necessary to comprehensively describe a hand while facilitating options to give rich details for various use cases beyond simple action tracking. For convenience in parser implementations, the formatting must also be simply structured in a consistent manner. In addition, the format must be powerful enough to describe a wide variety of poker games played in formal settings. Such selections should cover games played in reputable poker tournaments around the world while the inclusion of various eccentric or esoteric variants is not as necessary.
@@ V Specification
The PHH format is a derivative of the Tom’s Obvious, Minimal Language (TOML) format [17]. This design allows PHH files to take advantage of TOML format’s type system and be human-readable and writable while maintaining easy parsability by software systems. This differs from historical game file formats like PGN [10] or SGF [11] which use in-house formatting.
The PHH format puts restrictions on the naming and types of the key/values, with which the poker games are described. These definitions, named “fields,” fall under two classifications. The first involves state construction and progression and must be specified. The second describes various miscellaneous information about the game and may be omitted by the annotator.
@@ V-B Position
In typical button games such as no-limit Texas hold ’em or pot-limit Omaha hold ’em, the first player must be in the small blind (or big blind if heads-up) while the last player must be in position (i.e. has the button). In non-button games like stud, there is no strict notion of position. In such games, the first player should be the one in the dealer’s immediate left while the last player should be in the dealer’s immediate right.
@@ V-C Rules
As poker is a game with a diverse set of rules that often differ based on regions or events, it is impossible to point to anything as an authoritative documentation of poker rules that must be followed both by the annotator and the parser. The parser implementation should be designed to handle actions that may be potential transgressions in certain poker rule sets, either silently or with a warning.
@@ VI Required Fields
Variant code variant String
@@
Forced bets blinds_or_straddles Array of integers or floats
@@
Starting stacks starting_stacks Array of integers, floats, or null
@@
Figure 2: Example definitions in PHH format.
The required fields represent a handful of fields that the annotator must define to digitalize a poker hand and therefore form a valid PHH file. The required fields are either used during the initial state construction or the subsequent state progression.
In this paper, 11 variants are explored, each of which is written in an abbreviated form, as tabulated in Table I. These variants include all 9 played in the 2023 WSOP Event 43: \$50,000 PPC [12], in addition to 2 others like short-deck and badugi in which the poker community has shown growing interest. These were chosen as they adequately represent the diverse nature of poker variants and demonstrate the flexibility of the specification laid out in this paper which is powerful enough to accommodate all variants in the 2023 WSOP Tournament Rules [6]. Those not mentioned in Table I can simply be integrated by introducing new variant codes to be associated with, and they will be introduced on a need-by-need basis based on community feedback in the future version of the specification.
The fields “antes” and “blinds_or_straddles” represent the antes and forced bets (except the bring-in), respectively, posted by each of the players in the hand. Note that antes and blinds are reversed in heads-up games by rule. Example definitions are shown in Figure 2. The “blinds_or_straddles” are forced bets only relevant in button games while the “bring_in” is a type of forced bet only used in stud game variants.
The “small_bet” and “min_bet” fields denote the minimum bets in the first and last few betting rounds in fixed-limit games. In no-limit or pot-limit games, these two values are, by rule, identical and therefore combined into a single “min_bet” field.
The “starting_stacks” field denotes the starting stacks of each player and is an array of positive integers or floats of length equal to the number of players. All stack sizes must strictly be non-zero. Unknown stack values can be denoted as “null”.
TABLE IV: Action notations in PHH File Format.
Action Grammar
Dealing community cards d db card(s)[ # Commentary]
Dealing down/up cards d dh pn card(s)[ # Commentary]
Bringing in pn pb[ # Commentary]
Completing/Betting/Raising to pn cbr amount[ # Commentary]
Checking/Calling pn cc[ # Commentary]
Folding pn f[ # Commentary]
Standing pat/Discarding pn sd[ card(s)][ # Commentary]
Showing/Mucking their hole cards pn sm[ card(s)][ # Commentary]
No-ops [# Commentary]
The only field that represents the progression of a poker state is the “actions” field. This field is an array of string action representations, following the action notation described in table IV. It describes how the state mutates throughout the hand as the actors – the dealer and the players – act.
The notations in the “actions” field must be sorted in the order of the action. The actions may represent a complete history of the hand or a partial history that does not reach the terminal state (the hand being over).
@@
The standing pat and discarding actions are grouped together due to their similarity. Standing pat is a special case of discarding where a player does not discard any hole cards. This notation accepts an optional argument denoting the cards that are being discarded. An omission of this argument denotes that the player in turn is standing pat while the inclusion denotes that the player discarding.
The showdown action is performed at the end of the hand or after all players go all-in. It represents showing of the hole cards to try to win the pot or after an all-in, or mucking of the hole cards. This action accepts an optional argument of cards which, if supplied, must all be known. The omission of this field represents mucking while the inclusion represents showing.
VI-A Card Notation
TABLE V: All ranks and their single character representations.